### 15 Dec 2023 v2

# Mitigating Outlier Activations in Low-Precision Fine-Tuning of Language Models

Abstract
Low-precision fine-tuning of language models has gained prominence as a cost-effective and energy-efficient approach to deploying large-scale models in various applications. However, this approach is susceptible to the existence of outlier values in activation. The outlier values in the activation can negatively affect the performance of fine-tuning language models in the low-precision regime since they affect the scaling factor and thus make representing smaller values harder. This paper investigates techniques for mitigating outlier activation in low-precision integer fine-tuning of the language models. Our proposed novel approach enables us to represent the outlier activation values in 8-bit integers instead of floating-point (FP16) values. The benefit of using integers for outlier values is that it enables us to use operator tiling to avoid performing 16-bit integer matrix multiplication to address this problem effectively. We provide theoretical analysis and supporting experiments to demonstrate the effectiveness of our approach in improving the robustness and performance of low-precision fine-tuned language models.
1 INTRODUCTION
Language models have achieved remarkable success in various NLP tasks, owing to their ability to capture the intricacies of text data. Fine-tuning large language models, however, often requires substantial computational resources and memory bandwidth that hinder its accessibility to users with limited computational resources. To make large language models accessible and efficient for real-world applications, researchers have explored various techniques for making fine-tuning pre-trained models more efficient on devices with lower computational power. To mitigate these challenges, low-precision fine-tuning has emerged as a promising approach.
Low-precision fine-tuning involves representing the model’s weights, activations, and also gradients to lower bit-width representations, such as 8-bit integer or floating-point numbers. This approach reduces memory and computational requirements, making it feasible to fine-tune and deploy large-scale models on resource-constrained devices. However, both in fine-tuning and inference, this approach introduces the problem of outlier activation, where a small number of activations exhibit extreme values, causing numerical instability and degradation in model performance.
Figure 1: Computation flow of proposed linear layers for forward and backward propagation. Integer computation significantly reduces the computational cost of compute-intensive linear layers.
In this paper, we delve into the problem of outlier activation in low-precision fine-tuning of language models. In our proposed approach, weights, activations, and gradients of all compute-intensive linear layers are represented using integer number formats. Instead of quantization approaches used in the literature, we propose a comprehensive approach that uses various hardware design techniques to address this issue effectively. Our contributions are as follows.
  • We analyze the causes and consequences of outlier activation in low-precision fine-tuning. We find that outlier activations are more important in the forward pass. Thus, we keep all the gradients in the back-propagation of linear layers in 8-bit integers.
  • Instead of quantizing the floating point values, we switch the number format of weights, activations, and gradients to an adaptive-integer number format which considers different integer lengths (e.g. INT12 or INT16) for activation outliers (less than 5% of all parameters) and keeps all the other parameters in INT8 format.
  • Using the advantage of integer number formats, we present a tiling strategy that enables the possibility of using int8 GEMM for all the computation of linear layers. Note that such tiling strategy is not easily possible for floating-point number formats such as (FP16 and FP8)
  • We provide theoretical analysis on how treating outliers separately helps to preserve the information in low-precision regime.
In Figure 1, we present an overview of our novel linear layers, highlighting the innovative handling of outlier activations in an integer number format while maintaining gradient computation in low-precision integer format. Notably, to the best of our knowledge, this is the first fully INT8 linear layer designed to manage outlier features in integer format, while simultaneously preserving gradient calculations in low-precision integer format.
2 RELATED WORKS
The emergence of Large Language Models (LLMs) has revolutionized natural language processing, yet their formidable size poses significant computational challenges for training, fine-tuning, and deployment. To address these challenges, intensive research has focused on quantization techniques, low-precision arithmetic, and compression methods. This Section investigates these approaches and their efficacy in mitigating outlier activation during inference and back-propagation, offering insights into the evolving landscape of techniques designed to make LLMs more efficient and accessible in resource-constrained environments.
2.1 Handling Outliers in Low-precision Inference
Most of the research efforts in the literature are focused on studying the effect of outlier activations in the forward propagation i.e. inference. For instance, LLM.int8() proposed by [Dettmers et al., 2022] decompose the outlier activations and their corresponding weights to a separate matrix multiplication that is performed in FP16 format while keeping the values that are not outlier in INT8 format. They also show that using a threshold is enough for detecting the outlier features. GPTQ presented by [Frantar et al., 2022] is a one-shot post-training quantization (PTQ) scheme that is based on approximate second-order information. GPTQ quantizes the weights while keeping the activations in floating point format. AWQ proposed by [Lin et al., 2023] is another PTQ scheme that focuses on protecting salient activations by applying normalizing scales for weights and activation tensors. These scales are determined by only analyzing activation tensors. [Dettmers and Zettlemoyer, 2023] proposed an outlier-dependent quantization scheme called proxy quantization which quantizes the weights corresponding to the outliers into a higher precision number format. Proxy quantization exploits the standard deviation of each layer’s hidden unit weights as a proxy for which dimensions have outlier features. Outlier channel splitting (OCS) proposed by [Zhao et al., 2019] tackle the problem of outlier features by duplicating channels containing outliers, then halves the channel values. Morover, norm tweaking is proposed by [Li et al., 2023] to reverse the magnification of outliers by normalization layers i.e. LayerNorm as discovered by [Wei et al., 2022]. SmoothQuant proposed by [Xiao et al., 2023] offers an 8-bit quantization scheme for weights and activation. SmoothQuant deals with outlier features by migrating the quantization difficulty from activation to weights using scaling factors for weights and activations. Furthermore, [Dettmers et al., 2023] proposed SpQR that isolates outlier weights, which may cause large quantization errors, and stores them in higher precision, an then compresses all other weights to integer format. [Yuan et al., 2023] suggested RPTQ method which reorders the channels with outliers and group them in order to reduce the quantization error.
Figure 2: Inference operations in an integer-only linear layer. The bottom panel shows the linear fixed-point mapping for the input tensors, which can adapt different bit-widths for activation outliers and other parameters in the layer.
2.2 Handling Outliers in Low-precision Back-propagation
Over the past few years, low-precision training of deep learning models has gained popularity in reducing the training cost. For instance, FP16 mixed precision training [Micikevicius et al., 2017] is nowadays a commonplace methodology to fine-tune language models. Integer data type has also been extensively studied for back-propagation computations by [Zhang et al., 2020, Zhao et al., 2021, Zhu et al., 2020, Ghaffari et al., 2022]. Moreover, using higher-bit integer formats such as INT12 for both back-propagation and forward propagation is proposed by [Hosseini et al., 2023].
Nonetheless, the majority of literature pertinent to low-precision back-propagation and training does not address the emergence of outliers in language models. As a result, our paper is dedicated to investigating the impact of outliers in low-precision integer training of language models. In the remainder of the paper, we demonstrate the significance of outlier features in the forward pass. Additionally, we establish that the forward pass can be entirely computed using integer arithmetic. Finally, we emphasize that handling outlier separately is only important in the forward pass. As for the back-propagation, all the parameters can remain in INT8 format.
3 METHODOLOGY
This section delves into our proposed methodology for mitigating outlier values in the low-precision training of language models. We consider keeping all the parameters in the back-propagation in INT8 format while treating the outlier activation separately in the forward pass. We found that the outlier activation does not need to be treated differently (i.e. representing outliers in higher precision) in the back-propagation.
3.1 Number representation
We employ the dynamic fixed-point format, also referred to as block floating-point [Williamson, 1991], to convert floating-point numbers into integer data types. In this format, floating-point numbers are mapped into blocks of integer values, each assigned a unique scale.
To perform this conversion, we utilize a linear fixed-point mapping function, which transforms a floating-point tensor $\mathbf{F}$ into a tensor of integers along with a single scale factor. The integer values are derived by rounding the floating-point mantissas, while the scale is determined as the maximum of the floating-point exponents within $\mathbf{F}$. The operational details of the linear fixed-point mapping are illustrated in the lower section of Figure 2.
To convert the fixed-point integers back into floating-point representation, we employ a non-linear inverse mapping function. This inverse mapping function converts integer values into normalized floating-point mantissas, associating each integer with its respective scale before packaging them into a floating-point number.
For a more comprehensive insight into the representation mapping functions, readers can refer to [Ghaffari et al., 2022]. It is worth noting that our approach deviates from existing methods by introducing various bit-widths for outlier activations in the fine-tuning of transformer-based language models. This strategy enables us to explore different bit-width configurations for handling outlier activations, ultimately facilitating the determination of the minimum memory band-width required for fine-tuning language models both in forward and backward propagations.
3.2 Proposed methods
In this subsection, we present two approaches designed to mitigate the impact of outliers in the context of low-precision language model fine-tuning. The first approach, the unified scale for outliers, seeks to provide a consistent scaling mechanism for outlier activations. The second approach, splitting outlier activations (Tiling), explores a novel strategy to isolate and manage outlier activations effectively, taking advantage of having two scaling factors for outliers. Also note that in both approaches, we used a threshold $\gamma=5$ to isolate the outlier activations for all the linear layers.
3.2.1 Approach 1: Unified Scale for Outliers
In this approach, we completely isolate the outlier activations and their corresponding weights and quantize them to INT12 while the rest are quantized to INT8. Let us assume $\mathbf{X}$ denotes the activation tensor, and $Q()$ is the quantization function, then
$Q(\mathbf{X})=S_{\mathbf{x}}{\mathbf{X}}^{\texttt{INT8}}+S_{\text{outlier}}{\mathbf{X}}^{\texttt{INT12}}_{\text{outlier}}.$ (1)
3.2.2 Approach 2: Splitting Outlier Activations (Tiling)
In the second approach, the value of outliers are split into two values $\hat{\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP1}}$ and ${\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP2}}$ and the quantization scheme is as follows
$Q(\mathbf{X})=$ (2)
$S_{\mathbf{x}}({\mathbf{X}}^{\texttt{INT8}}+{\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP1}})+S_{\text{outlier}}{\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP2}}.$
In this quantization scheme, we extract floating-point outlier activations $\mathbf{X}_{\text{outlier}}$ from original floating-point activations $\mathbf{X}$ using threshold $\gamma$ and then, split them as shown in the following equations,
$\mathbf{X}_{\text{outlier\_SP2}}=\lfloor\frac{\mathbf{X}_{\text{outlier}}+\gamma}{2\gamma}\rfloor\times 2\gamma$ (3)
$\mathbf{X}_{\text{outlier\_SP1}}=\mathbf{X}_{\text{outlier}}-\mathbf{X}_{\text{outlier\_SP2}}.$
and then we quantize them to get ${\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP1}}$ and ${\mathbf{X}}^{\texttt{INT8}}_{\text{outlier\_SP2}}$. The benefit of this method is that we can keep the computation of forward pass completely in INT8 format while treating the outlier separately from the no-outlier activation values.
4 THEORETICAL ANALYSIS
In this section, we delve into the implications of low-precision number formats on information preservation. The utilization of reduced bit-width representations in deep learning, while advantageous for efficiency and resource conservation, inevitably introduces the issue of information loss. We explore the nuances of this phenomenon and employ sensitivity analysis to quantify the extent to which information is altered or discarded in the transition from high precision to low precision.
Furthermore, we extend our investigation to consider distribution distances, such as the $\chi^{2}$-divergence and the Hammersley–Chapman–Robbins bound.
4.1 Information Loss in Low-precision Number Formats
The concept of sample informativeness is a well-established notion within the field of statistics. For example, [Tukey, 1965] introduced a dimensionless metric for measuring informativeness, which proves particularly valuable for our analysis. To measure the informativeness, [Tukey, 1965] defines the concept of leverage and linear sensitivity as
$\mathrm{lev}_{\theta}(X)=\frac{\partial}{\partial\theta}\mathbb{E}_{\theta}(X)$
$\mathrm{sens}_{\theta}(X)=\frac{(\mathrm{lev}_{\theta}(X))^{2}}{\mathbb{V}(X)}.$ (4)
Thus, we can re-write the equation (4) in the low-precision number formats $\hat{X}$ if we consider a low precision number has a rounding error of $\delta$ in a way that $\hat{X}=X+\delta$. Note that we assume $\delta$ and $X$ are independent random variables and $\mathbb{E}(\delta)=\varepsilon\simeq 0$.
$\frac{\mathrm{lev}_{\theta}(\hat{X})}{\mathrm{lev}_{\theta}(X)}\simeq 1~{}~{}~{}~{}~{}~{}\text{s.t}~{}~{}~{}~{}\mathbb{E}(\delta)\simeq 0$
$\frac{\mathrm{sens}_{\theta}(\hat{X})}{\mathrm{sens}_{\theta}(X)}=\frac{\mathbb{V}(X)}{\mathbb{V}(\hat{X})}=\frac{\mathbb{V}(X)}{\mathbb{V}(X)+\mathbb{V}(\delta)}\leq 1.$ (5)
Inequality (5) shows that in the case of unbiased rounding, low-precision representation always increases the variance and therefore decreases the informativeness of the sample.
It is noteworthy to mention that the sensitivity measure defined in equation (4) is closely related to Hammersley-Chapman-Robbins lower-bound [Chapman and Robbins, 1951],
$\frac{\left(\ \mathbb{E}(X)-\mathbb{E}(\hat{X})\right)^{2}}{\mathbb{V}(\hat{X})}\leq$
$\frac{\left(\mathbb{E}(X)-\mathbb{E}(\hat{X})\right)^{2}}{\mathbb{V}(X)}\leq\chi^{2}(f_{X}||f_{\hat{X}}),$ (6)
which provides a lower-bound for $\chi^{2}$-divergence of $X$ and $\hat{X}$ distributions. Note that $\chi^{2}$-divergence is a measure to quantify the divergence between two distributions and for distributions $P$ and $Q$ is defined as,
$\chi^{2}(P||Q)=\int(\frac{dP}{dQ}-1)^{2}dQ.$ (7)
Figure 3: Outliers modeled as a mixture distribution.
4.2 Analysing Outlier Activations as a Mixture Distribution
Treating outlier activations separately as explained in Section 3.2 closely resembles having a mixture distribution as shown in Figure 3. This means that we consider the outlier activations are samples that are drawn from a different distribution function than non-outlier activations. Let us assume the original distribution of $X$ as $f_{X}$ and a threshold $\gamma$ that separates outliers $f_{X_{2}}$ from the rest of activations $f_{X_{1}}$. define $F_{X}$ as the coefficients of $X$ such that
$f_{X}(x)=pf_{X_{1}}(x)+(1-p)f_{X_{2}}(x),$
$p=F_{X}(\gamma),$
$f_{X_{1}}(x)=\frac{f_{X}(x)\mathcal{I}_{\{X\leq\gamma\}}}{F_{X}(\gamma)},$
$f_{X_{2}}(x)=\frac{f_{X}(x)\mathcal{I}_{\{X>\gamma\}}}{1-F_{X}(\gamma)}.$ (8)
Let us further assume $Y=\mathcal{I}_{\{X\leq\gamma\}}$, then
$\mathbb{E}(X)=\mathbb{E}\left(\mathbb{E}(X|Y)\right)$
$=p\mathbb{E}_{f_{X_{1}}}(X)+(1-p)\mathbb{E}_{f_{X_{2}}}(X),$ (9)
and,
$\mathbb{E}(X^{2})=p\mathbb{E}_{f_{X_{1}}}(X^{2})+(1-p)\mathbb{E}_{f_{X_{2}}}(X^{2}).$ (10)
Furthermore, by subtracting the $p\mathbb{E}^{2}_{f_{X_{1}}}(X)+(1-p)\mathbb{E}^{2}_{f_{X_{2}}}(X)$ from (10) and using Jensen’s inequality we have
$\mathbb{V}(X)\geq p\mathbb{V}_{f_{X_{1}}}(X)+(1-p)\mathbb{V}_{f_{X_{2}}}(X).$ (11)
Remark 1. The inequality (11) shows that weighted average of variances of distibutions is less than total variance of a mixture distribution. Therefore, treating outlier separately reduces the variance and hence it increases the informativeness i.e. sensitivity according to equation (4).
STSB QNLI MNLI SST-2 RTE MRPC CoLA
FP32 87.6 89.9 83.5 91.9 61.7 78.7/85.3 55.3
[…]
Proposed Approach 2 81.6 89.6 82.6 91.5 59.2 74.3/83.5 52.2
INT8 Untreated Outliers 80.5 86.4 80.9 91.8 58.5 69.9/81.9 47.0
Table 1: Metric performance of integer fine-tuning of BERT on selected GLUE tasks. The reported metric for QQP and MRPC is accuracy and F1 score, for QNLI, MNLI, RTE, and SST-2 is accuracy, for STSB is the Pearson-Spearman correlation, and for CoLA is the Matthews correlation.
SQuAD v1.1 SQuAD v2
FP32 79.6/87.5 71.5/74.8
[…]
Proposed Approach 2 74.9/84.1 65.5/69.0
INT8 Untreated Outliers 69.8/80.2 60.9/64.6
Table 2: Metric performance of fine-tuning BERT on SQuAD v1.1 and v2.0 datasets. For both datasets, the exact match metrics and F1 scores are reported.
4.3 Informativeness of Mixture Distribution in Low-Precision Number Formats
In this section, we try to re-establish the results of Section 4.2 for low-precision number formats. To do so, we need to show equation (8) holds in the low-precision number format.
Let us consider the following low-precision representations, $\hat{X}=X+\delta$, $\hat{X}_{1}=X_{1}+\delta$ and $\hat{X}_{2}=X_{2}+\delta$, where $\delta$ is the rounding error and is independent from $X$. The moment generating function $m_{\hat{X}}$ is
$m_{\hat{X}}(t)=\mathbb{E}(e^{t\hat{X}})=\mathbb{E}(e^{tX})\mathbb{E}(e^{t\delta})$
$=m_{X}(t)m_{\delta}(t).$ (12)
Now,
$m_{X}(t)$ $=\int_{-\infty}^{\infty}e^{tX}f_{X}(x)$
$=p\int_{-\infty}^{\infty}e^{tX_{1}}f_{X_{1}}(x)$
$+(1-p)\int_{-\infty}^{\infty}e^{tX_{2}}f_{X_{2}}(x)$
$=pm_{X_{1}}(t)+(1-p)m_{X_{2}}(t).$ (13)
and thus, using equations (12) and (13),
$m_{\hat{X}}(t)$ $=m_{X}(t)m_{\delta}(t)$
$=pm_{X_{1}}(t)m_{\delta}(t)+(1-p)m_{X_{2}}(t)m_{\delta}(t)$
$=pm_{\hat{X}_{1}}(t)+(1-p)m_{\hat{X}_{2}}(t).$ (14)
which means,
$f_{\hat{X}}(x)=pf_{\hat{X}_{1}}(x)+(1-p)f_{\hat{X}_{2}}(x).$ (15)
Remark 2. Establishing equation (15) confirms that inequality (11) holds for the low-precision regime and thus, treating outlier activations separately in low-precision reduces the quantization variance (i.e. quantization noise) and increases the informativeness.
Remark 3. The equation (15) holds if the moement generative functions exist. This is a valid assumption since the distribution of activation has bounded support.
5 EXPERIMENTAL RESULTS
5.1 Experiment Setup
We conducted fine-tuning on the BERT base model across a series of downstream tasks to facilitate a performance comparison between our integer fine-tuning method and the FP16 and FP32 fine-tuning approaches. The fine-tuning process encompassed specific tasks selected from the GLUE benchmark [Wang et al., 2018], in addition to the Stanford Question Answering Datasets, specifically SQuAD v1.1 and SQuAD v2.0 [Rajpurkar et al., 2016].
Each fine-tuning setup was standardized with identical hyper-parameters and an equivalent number of training epochs. To ensure result stability, reported metrics represent the average of five runs, each initialized with a different random seed to mitigate the impact of random variations.
Our fine-tuning experiments were executed using the fine-tuning scripts provided by the Hugging Face library [Wolf et al., 2019]. In the case of GLUE experiments, fine-tuning spanned five epochs, with a learning rate set to $2\times 10^{-5}$, and a per-device fine-tuning batch size of 32. Meanwhile, the fine-tuning of BERT on the SQuAD datasets comprised two epochs, with a learning rate of $5\times 10^{-5}$ and a per-device fine-tuning batch size of 12. Notably, all experiments were conducted on a computational infrastructure consisting of eight NVIDIA V100 GPUs, each equipped with 32 gigabytes of VRAM.
5.2 Results
The results of fine-tuning BERT base on GLUE benchmark and SQuAD datasets are presented in Table 1 and Table 2 respectively. Our proposed solution significantly improves the robustness of low-precision fine-tuning for BERT on GLUE and SQuAD datasets. Comparing the results of the proposed approaches with INT8 fine-tuning with untreated outliers shows that representing outliers separately almost always improves the fine-tuning performance of the model. Additionally, we emphasize again that, the results presented in this paper comprise of having both back-propagation and forward-propagation in low-precision number formats and show that low-precision arithmetic are promising avenue to reduce the computational complexity of language models.
6 CONCLUSION
This paper explored means to mitigate the outlier activations in low-precision language model fine-tuning. We have introduced a novel methodology for mitigating the challenges posed by outlier activations, offering effective approaches as effective approaches to enhance the stability of the fine-tuning phase in low precision number format where gradients, weights, and activations are in INT8 format. Additionally, we provided a theoretical analysis to understand the intricacies of information loss in low-precision number formats. Our sensitivity analysis has unveiled the trade-offs between variance and informativeness while considering distribution distances like the $\chi^{2}$-divergence and the Hammersley–Chapman–Robbins bound has deepened our insights into these transformations. In a landscape where the deployment of large language models is increasingly resource-constrained, our work contributes to the ongoing efforts to make these models more accessible and efficient. By addressing the challenges of outliers and information loss, we pave the way for the continued evolution of low-precision back-propagation in language model fine-tuning. Our findings not only have implications for natural language processing but also hold relevance for broader applications across data analysis and machine learning.

### 13 Jan 2024 v3

@@
 Integer data type has also been extensively studied for back-propagation computations by [Zhang et al., 2020, Zhao et al., 2021, Zhu et al., 2020, Ghaffari et al., 2022].
-Moreover, using higher-bit integer formats such as INT12 for both back-propagation and forward propagation is proposed by [Hosseini et al., 2023].
+Moreover, using higher-bit integer formats such as INT12 for both back-propagation and forward propagation is proposed by [Tayaranian et al., 2023].
 Nonetheless, the majority of literature pertinent to low-precision back-propagation and training does not address the emergence of outliers in language models.
@@
 $\mathrm{lev}_{\theta}(X)=\frac{\partial}{\partial\theta}\mathbb{E}_{\theta}(X)$
-$\mathrm{sens}_{\theta}(X)=\frac{(\mathrm{lev}_{\theta}(X))^{2}}{\mathbb{V}(X)}.$ (4)
+$\mathrm{sens}_{\theta}(X)=\frac{(\mathrm{lev}_{\theta}(X))^{2}}{\mathbb{V}(X)},$ (4)
+where $\theta$ is the parameters of $X$ distribution.
 Thus, we can re-write the equation (4) in the low-precision number formats $\hat{X}$ if we consider a low precision number has a rounding error of $\delta$ in a way that $\hat{X}=X+\delta$.
@@
 sensitivity according to equation (4).
+Table 1: Metric performance of integer fine-tuning of BERT on selected GLUE tasks.
+The reported metric for MRPC is accuracy and F1 score, for QNLI, MNLI, RTE, and SST-2 is accuracy, for STSB is the Pearson-Spearman correlation, and for CoLA is the Matthews correlation.
 STSB QNLI MNLI SST-2 RTE MRPC CoLA
 FP32 87.6 89.9 83.5 91.9 61.7 78.7/85.3 55.3
+FP16 88.6 90.1 83.2 91.7 59.6 77.7/85.1 56.0
 […]
 Proposed Approach 2 81.6 89.6 82.6 91.5 59.2 74.3/83.5 52.2
-INT8 Untreated Outliers 80.5 86.4 80.9 91.8 58.5 69.9/81.9 47.0
-Table 1: Metric performance of integer fine-tuning of BERT on selected GLUE tasks.
-The reported metric for QQP and MRPC is accuracy and F1 score, for QNLI, MNLI, RTE, and SST-2 is accuracy, for STSB is the Pearson-Spearman correlation, and for CoLA is the Matthews correlation.
+INT8 Untreated Outliers 80.9 86.4 80.9 91.8 58.5 69.9/81.9 43.5
+Table 2: Metric performance of fine-tuning BERT on SQuAD v1.1 and v2.0 datasets. For both datasets, the exact match metrics and F1 scores are reported.
 SQuAD v1.1 SQuAD v2
 FP32 79.6/87.5 71.5/74.8
+FP16 79.6/87.5 69.1/72.2
 […]
@@
 INT8 Untreated Outliers 69.8/80.2 60.9/64.6
-Table 2: Metric performance of fine-tuning BERT on SQuAD v1.1 and v2.0 datasets. For both datasets, the exact match metrics and F1 scores are reported.
 4.3 Informativeness of Mixture Distribution in Low-Precision Number Formats